# Study on data analysis for Ives-Stilwell-type experiments

### v14

@@ Abstract
Ives-Stilwell experiment in 1938 is a historic experiment for confirming Einstein’s special relativity, and various modern types have been repeated by use of laser technology. However in this paper, we reveal and solve a fundamental issue that the data analysis for all those experiments is not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect and its associated time dilation have not actually been confirmed. For example, in the Letter [Phys. Rev. Lett. 113, 120405 (2014)] the definition of the measurement accuracy of Doppler effect, given by $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$, is not physical because Einstein’s Doppler formula cannot be confirmed even when $\varepsilon=0$ holds (see Appendix A). We argue that there are two first principles for analyzing and confirming Einstein’s Doppler effect, stating: (i) Einstein’s Doppler effect refers to the same photon (or laser beam) exhibiting different frequencies observed in different inertial frames, and (ii) the quantity (or measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. Unfortunately, the data analysis for the 1938 Ives-Stilwell experiment does not comply with the first principles, so it fails to confirm the relativistic effect (see Appendix B), although this data analysis has been mimicked by quite a few generations of physicists. (It should be emphasized that in fact, Ives-Stilwell data analysis and the experimental test itself both fail to confirm Einstein’s Doppler effect.) Based on the first principles, we propose a justified data analysis and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus resulting in a great advance in the experimental verification of Lorentz invariance via the Doppler effect.
pacs:
03.30.+p, 03.50.De
@@ I Introduction
In this paper, we argue that there are two first principles for analyzing and confirming Einstein’s Doppler effect, stating: (i) Einstein’s Doppler effect refers to the same photon (or laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (or measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. Unfortunately, the way of thinking set up in the 1938 experiment did not comply with the first principles so that even Ives and Stilwell themselves were unaware that their experimental test actually did not support Einstein’s Doppler effect (see Appendix B). Here, based on the above first principles with existing experimental data employed r7 , we propose a justified analysis method and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus successfully revealing and solving a long-lasting unrecognized fundamental issue in the experimental confirmation of special relativity.
@@ Appendix B Ives-Stilwell experimental test failing to confirm Einstein’s Doppler effect
where $\lambda_{0}$ is the wavelength observed in the moving frame, such as in the ion-rest frame in the Ives-Stilwell case, $\lambda_{\textrm{lab}}$ is the wavelength observed in the laboratory frame, $\boldsymbol{\beta}$ is the velocity at which the frame moves relatively to the laboratory frame, and $\hat{\mathbf{n}}$ is the unit wave vector. From this we can see that the relativistic Doppler effect exists observed in any directions in the laboratory fame as long as $\boldsymbol{\beta}\neq 0$ or $\gamma=(1-\boldsymbol{\beta}^{2})^{-1/2}>1$ holds. However in the Ives-Stilwell experiment test, out of all 8 cases, each has a light beam that does not have the relativistic effect. For example, for case 8 with $D_{b}=-0.1317<0$ leading to $\varepsilon_{b}=-3.7\times 10^{-5}<\varepsilon_{c}=-9.8\times 10^{-6}$, the backward beam has a measured wavelength of 4882.264882.26Å less than its classic Doppler wavelength of 4882.394882.39Å, without any relativistic effect.
@@ Acknowledgments
Data underlying the results presented in this paper are available within the paper and Supplement 1 r13 .

### v15

@@ Abstract
Ives-Stilwell experiment in 1938 is a historic experiment for confirming Einstein’s special relativity, and various modern types have been repeated by use of laser technology. However in this paper, we reveal and solve a fundamental issue that the data analysis for all those experiments is not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect and its associated time dilation have not actually been confirmed. We argue that there are two first principles for analyzing and confirming Einstein’s Doppler effect, stating: (i) Einstein’s Doppler effect refers to the same photon (or laser beam) exhibiting different frequencies observed in different inertial frames, and (ii) the quantity (or measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. Unfortunately, Ives-Stilwell data analysis method does not comply with the first principles, failing to find that the experimental data they provided actually does not support Einstein’s Doppler effect. Robertson-Mansouri-Sexl test theory is widely used to test Lorentz invariance, but it does not adhere to the first principles either when employed to measure time dilation. Based on the first principles, we propose a justified data analysis and correctly confirm the Doppler effect and its associated time dilation in the Ives-Stilwell-type experiment.
@@ I Introduction
In this paper, we argue that there are two first principles for analyzing and confirming Einstein’s Doppler effect, stating: (i) Einstein’s Doppler effect refers to the same photon (or laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (or measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. Unfortunately, Ives and Stilwell performed a data analysis for their experiment test r2 that did not comply with the first principles so that they even failed to find that the experimental data they provided actually did not support relativistic effect (see Appendix B). Robertson-Mansouri-Sexl (RMS) test theory is widely used to test Lorentz invariance; however, this RMS theory does not adhere to the first principles either when employed to measure time dilation in Ives-Stilwell-type experiments r6 (see Appendix C). Here, based on the above first principles with existing experimental data employed r7 , we propose a justified analysis method and correctly confirm the Doppler effect and its associated time dilation in the Ives-Stilwell-type experiment, thus successfully revealing and solving a long-lasting unrecognized fundamental issue in the experimental confirmation of special relativity.
@@ Appendix B Ives-Stilwell experimental test failing to confirm Einstein’s Doppler effect
where $\lambda_{0}$ is the wavelength observed in the moving frame, such as in the ion-rest frame in the Ives-Stilwell case, $\lambda_{\textrm{lab}}$ is the wavelength observed in the laboratory frame, $\boldsymbol{\beta}$ is the velocity at which the frame moves relatively to the laboratory frame, and $\hat{\mathbf{n}}$ is the unit wave vector. From this we can see that the relativistic Doppler effect exists observed in any directions in the laboratory frame as long as $\boldsymbol{\beta}\neq 0$ or $\gamma=(1-\boldsymbol{\beta}^{2})^{-1/2}>1$ holds. However in the Ives-Stilwell experiment test, out of all 8 cases, each has a light beam that does not have the relativistic effect. For example, for case 8 with $D_{b}=-0.1317<0$ leading to $\varepsilon_{b}=-3.7\times 10^{-5}<\varepsilon_{c}=-9.8\times 10^{-6}$, the backward beam has a measured wavelength of 4882.264882.26Å less than its classic Doppler wavelength of 4882.394882.39Å, without any relativistic effect. (See Ref. r15 for all detailed numerical analysis of the Ives-Stilwell experimental test.)
@@
Appendix C Time dilation failed to be confirmed in the optical-atomic-clock experiment
Reinhardt and coworkers r6 reported that time dilation is confirmed to unprecedented precision with a Mansouri-Sexl parameter $|\alpha|\leq 8.4\times 10^{-8}$ in an Ives-Stilwell-type experiment based on fast optical atomic clocks. However, we would like to point out that the data analysis method for this experiment is flawed, and thus the conclusion obtained in the paper r6 is under question, which is shown below.
In this paper r6 , a test theory by Robertson, Mansouri and Sexl (RMS) is adopted to analyze their experiment, where RMS transformation is a generalized Lorentz transformation, and the RMS transformation of the Ives-Stilwell experiment is modified as
$\frac{\nu_{a}\nu_{p}}{\nu_{0}^{2}}=1+2\alpha\beta^{2},$ (43)
where $\alpha$ is the Mansouri-Sexl parameter; $\nu_{a}$ and $\nu_{p}$ are, respectively, the frequencies of two laser beams; $\nu_{0}$ is the transition frequency of ions; $\beta$ is the ion’s moving velocity.
According to this theory, the time dilation will be confirmed when
$\alpha=0$ (44)
holds. Unfortunately, the RMS theory is not applicable in such a case, unless the authors, Reinhardt and coworkers r6 , assume that $\nu_{a}\nu_{p}=\nu_{0}^{2}$ only comes from Einstein’s Doppler formula
$\nu_{p,a}=\gamma(1\pm\beta)\nu_{0},$ (45)
as done in r13 , or that the relation $\nu_{a}\nu_{p}=\nu_{0}^{2}$ is valid only in special relativity, as claimed by Lämmerzahl r14 . In fact, however, even if the special relativity does not hold or Lorentz invariance is violated,
$\nu_{p,a}=10^{\pm 100}\gamma(1\pm\beta)\nu_{0}$ (46)
for example, we also can get the $\beta$-independent relation $\nu_{a}\nu_{p}=\nu_{0}^{2}$. Thus even if $\nu_{a}\nu_{p}=\nu_{0}^{2}$ or $\alpha=0$ exactly holds, Einstein’s Doppler formula or time dilation cannot be confirmed.
The failure of RMS test theory in this paper r6 lies in the fact that the Mansouri-Sexl parameter $\alpha$ is designed to measure time dilation via the combined Doppler effect from two different laser beams, instead of the effect from a single laser beam, which is in violation of first principles: (i) Einstein’s Doppler effect refers to the same photon exhibiting different frequencies observed in different inertial frames, and (ii) the quantity used as a measure to confirm the Doppler effect must be able to confirm Einstein’s Doppler formula itself.
@@ Acknowledgments
Data underlying the results presented in this paper are available within the paper and Supplement 1 r15 .